OpenStudy (anonymous):

3. An electron orbits the nucleus at a distance of 5.3x10-11 m and makes 6.6x1015 revolutions per second. find centripetal force

OpenStudy (anonymous):

@Loser66 i am a bit confused with this one

OpenStudy (loser66):

what do you get so far?

OpenStudy (anonymous):

i think the radius is 5.3 x 10^-11 and i understand what the 6.6x 10 ^15 revolution per second is?

OpenStudy (anonymous):

i dont have mass either

OpenStudy (anonymous):

i dont know, i think 1 second?

OpenStudy (anonymous):

so t is( 6.6 x 10 ^15) / 1 second

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I dont know what to do with this, I think i have time and i know the radius but i dont have mass.

OpenStudy (anonymous):

would i use a constant for mass of an electron

ganeshie8 (ganeshie8):

a = w^2R F = ma

OpenStudy (anonymous):

dont we use centripetal force formulas?

OpenStudy (anonymous):

we dont have w either?

ganeshie8 (ganeshie8):

use watever u have

ganeshie8 (ganeshie8):

or knw :)

OpenStudy (anonymous):

would we use w = 2 (pi) / r f = m r w^2

OpenStudy (anonymous):

i mean w = 2 (pi) r / t

ganeshie8 (ganeshie8):

wat is centripetal force formula ?

OpenStudy (anonymous):

f = m v^2 / r

ganeshie8 (ganeshie8):

or f = m w^2 R

ganeshie8 (ganeshie8):

w = angular velocity

ganeshie8 (ganeshie8):

anything wil do use watever u r at peace wid :)

OpenStudy (anonymous):

how do i find w

ganeshie8 (ganeshie8):

w = 2 pi f = 2*pi*6.6x10^15

OpenStudy (anonymous):

how would that be times by i thought we divide, I still dont understand the 6.6x1015 revolutions per second.

ganeshie8 (ganeshie8):

ever heard of the term 'frequency' beofore ?

OpenStudy (anonymous):

a little not too much in chapters

ganeshie8 (ganeshie8):

@Loser66 interpretation looks simple, see if u get it

ganeshie8 (ganeshie8):

i liked ur interpretation relating frequency directly to velocity :)

ganeshie8 (ganeshie8):

btw, thats angular velocity w. wat we talking so far

OpenStudy (anonymous):

okay so I can use w = 2 (pi) / r f = m r w^2 r is given the 6.6... mass is constant and w is 2pi x (6.6...)

ganeshie8 (ganeshie8):

w = 2*pi*f = 2*pi*6.6x10^15 r = 5.3x10-11 m = you knw it F = m w^2 R

OpenStudy (anonymous):

okay thank you

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

i am getting an answer of 8.30 x10^-8

ganeshie8 (ganeshie8):

wolfram says 65794 ?

OpenStudy (anonymous):

i dont know what i am doing wrong, I keep getting that answer

ganeshie8 (ganeshie8):

lay back... relax and work it again after some time :)

OpenStudy (anonymous):

w = 2 pi (6.6x10^15) = 4.146 x 10 ^16 f = (9.1 x 10^-31) (5.3 x 10^-11) (4.146 x 10 ^16)^2 = 8.3 x 10^-8

OpenStudy (anonymous):

i am putting ( ) around everything and it still doesnt work...i dont think wolf.. is correct ?

OpenStudy (anonymous):

someone else got the same answer as me too =(

ganeshie8 (ganeshie8):

Oh ya ! you're right !!! ive plugged it incorrectly. here is the correct link :- http://www.wolframalpha.com/input/?i=w+%3D+2*pi*6.6*10%5E15%2C+r+%3D+5.3*10%5E%28-11%29%2C+m%3D9.10938291*10%5E%28-31%29%2C+m*w%5E2*r+%3D+%3F

OpenStudy (anonymous):

omg thank you, i thought i was going insanee

ganeshie8 (ganeshie8):

good you're alert enough to catch my mistake =))

OpenStudy (anonymous):

thanks =)

ganeshie8 (ganeshie8):

np :D